# In-Context Learning Strategies Emerge Rationally

### v1

@@ H.1.2 Linear Regression
Figure 15: Task Diversity Threshold across Linear Regression conditions. Red dashed line indicates the memorizing solution $M$, blue dashed line indicates the generalizing solution $G$, and black solid line indicates Transformer behavior at the end of training ($100$K steps).
@@ H.1.3 Classification
Figure 16: Task Diversity Threshold across Classification conditions. Red dashed line indicates the memorizing solution $M$, blue dashed line indicates the generalizing solution $G$, and black solid line indicates Transformer behavior at the end of training ($100$K steps). IWL evaluation presented.
@@ H.2.3 Classification
Figure 19: Transience across Classification conditions with 8 dimensions. OOD performance presented. Blue Dashed line indicates OOD performance of generalizing solution $G$. Note that in the case of classification, it is often the case that only one or two task diversity conditions show transient generalization, as can be seen more clearly from the absolute distance maps in the next section.
@@ Appendix I Functional Form Ablations
Figure 27: Power laws over simplicity bias and sample efficiency are necessarily for explaining ICL phenomenology. By ablating our functional form, we see that the free parameters $\alpha,\beta,\gamma$ are required for the performance of the model. In particular, the simplicity bias derived from $K$ without a $\beta$ term is much too sharp and over-penalizes complexity compared to the Transformer. Additionally, memorization proceeds much too rapidly without the $\alpha$ term, pointing toward the necessity of assuming sub-linear sample efficiency for capturing Transformer training dynamics.
Appendix J Memorization Continues to Increase After Task Diversity Threshold - Refutation of Raventós et al. [25]’s Claim
@@
In our main experimental settings, in which we train with a constant learning rate, we find that relative distance trajectories follow a sigmoidal growth pattern with respect to $N^{1-\alpha}$ (see Fig. 6 and Fig. 29(a) top and bottom right). However, in contrast with our theory’s predicted sigmoidal curves which plateau at $1$ (i.e., some amount of training would yield full adherence to the memorizing solution, when it has a lower loss than the generalizing solution), this does not seem to be the case with the trajectories displayed by Transformers, which appear to plateau early (see Fig. 29(a) top right $D=256$ for a clear example). Indeed, fitting parameterized logistic curves to these trajectories yields plateau values different from $1$. To explain this, we turn to foundational work from Geman and Geman [87], who showed that a slow (logarithmic) temperature cooling schedule for Gibbs sampling substantially increases the likelihood of convergence to the global minimum (MAP estimate). Drawing a very rough parallel to the case of deep learning, it is reasonable to assume that a learning rate annealing schedule of some form is required to converge to the MAP estimate (which is the memorizing solution, in cases where it has lower loss than the generalizing solution). To test this, we repeated experiments in the Balls & Urns and Linear Regression settings and used warm-up and inverse-squared learning rate decay. Surprisingly, we indeed find that adding learning rate annealing can increase adherence to Bayes-optimal trajectories (Fig. 29(a)). However, we also find this effect is highly sensitive to training conditions: training for longer, even in conditions that yield the effect for a smaller number of training steps, can lead to plateau (Fig. 29(b), top), and slight changes in training conditions, in this case number of warm-up steps, can substantially reduce the effect (Fig. 29(b), bottom). It should also be noted that in this case, adherence to the Bayes-optimal trajectory is actually negative for generalization, since it means the model will converge to the memorizing solution quicker. However, this is not necessarily the case in more realistic training regimes, where the ability to overcome a simplicity bias and adopt more complex solutions is likely beneficial.

### v2

@@ H.1.2 Linear Regression
Figure 15: Task Diversity Effects Across Linear Regression Conditions. Red dashed line indicates the memorizing solution $M$, blue dashed line indicates the generalizing solution $G$, and black solid line indicates Transformer behavior at the end of training ($100$K steps).
@@ H.1.3 Classification
Figure 16: Task Diversity Effects Across Classification Conditions. Red dashed line indicates the memorizing solution $M$, blue dashed line indicates the generalizing solution $G$, and black solid line indicates Transformer behavior at the end of training ($100$K steps). IWL evaluation presented.
@@ H.2.3 Classification
Figure 19: Transience Across Classification Conditions. OOD performance presented. Blue Dashed line indicates OOD performance of generalizing solution $G$. Note that in the case of classification, it is often the case that only one or two task diversity conditions show transient generalization, as can be seen more clearly from the absolute distance maps in the next section.
@@ Appendix I Functional Form Ablations
Figure 27: Power laws over complexity measure and sample efficiency are necessarily for explaining ICL phenomenology. By ablating our functional form, we see that the free parameters $\alpha,\beta,\gamma$ are required for the performance of the model. In particular, the simplicity bias derived from $K$ without a $\beta$ term is much too sharp and over-penalizes complexity compared to the Transformer. Additionally, memorization proceeds much too rapidly without the $\alpha$ term, pointing toward the necessity of assuming sub-linear sample efficiency for capturing Transformer training dynamics.
Appendix J Memorization Continues to Increase After Task Diversity Threshold – Refutation of Raventós et al. [25]’s Claim
@@
In our main experimental settings, in which we train with a constant learning rate, we find that relative distance trajectories follow a sigmoidal growth pattern with respect to $N^{1-\alpha}$ (see Fig. 6 and Fig. 29(a) top and bottom right). However, in contrast with our theory’s predicted sigmoidal curves which plateau at $1$ (i.e., some amount of training would eventually yield full adherence to the memorizing solution, when it has a lower loss than the generalizing solution), this does not seem to be the case with the trajectories displayed by Transformers, which appear to plateau early (see Fig. 29(a) top right $D=256$ for a clear example). Indeed, fitting parameterized logistic curves to these trajectories yields plateau values different from $1$. To explain this, we turn to foundational work from Geman and Geman [98], who showed that a slow (logarithmic) temperature cooling schedule for Gibbs sampling substantially increases the likelihood of convergence to the global minimum (MAP estimate). Drawing a very rough parallel to the case of deep learning, it is reasonable to assume that a learning rate annealing schedule of some form is required to converge to the MAP estimate (which is the memorizing solution, in cases where it has lower loss than the generalizing solution). To test this, we repeated experiments in the Balls & Urns and Linear Regression settings and used warm-up and inverse-squared learning rate decay. Surprisingly, we indeed find that adding learning rate annealing can increase adherence to Bayes-optimal trajectories (Fig. 29(a)). However, we also find this effect is highly sensitive to training conditions: training for longer, even in conditions that yield the effect for a smaller number of training steps, can lead to plateau (Fig. 29(b), top), and slight changes in training conditions, in this case number of warm-up steps, can substantially reduce the effect (Fig. 29(b), bottom). It should also be noted that in this case, adherence to the Bayes-optimal trajectory is actually negative for generalization, since it means the model will converge to the memorizing solution quicker. However, this is not necessarily the case in more realistic training regimes, where the ability to overcome a simplicity bias and adopt more complex solutions is likely beneficial.